OpenStudy (anonymous):

Given the balanced equation 2 C4H10 (g) + 13 O2(g) = 8 CO2(g) +10 H2O(g) What is the total number of moles of O2(g) that must react completely with 1.54 moles of C4H10(g)?

OpenStudy (ipwnbunnies):

We're having fun, right? Use stoichiometry to find he mole ratio of C4H10 to O2, then multiply it by the amount of moles of C4H10 you got.

OpenStudy (ipwnbunnies):

I meant mole ratio of O2 to C4H10. We want to cancel out the moles of C4H10 in out math.

OpenStudy (anonymous):

lol, Thanks for helping. I only have 10 questions off of a 200 question review. I got 20.0 when I did that. Is that right?

OpenStudy (ipwnbunnies):

No, almost. You probably forgot the other part of the ratio, dividing by the C4H10 part.

OpenStudy (ipwnbunnies):

\[(1.54 mol C4H10)*(\frac{13 mol O2}{2 mol C4H10})\]

OpenStudy (anonymous):

umm.. when I did that I got 2 different answers... 1. 10.0 and 2. 26.5

OpenStudy (ipwnbunnies):

Hmm, make sure you put it in your calculator right.

OpenStudy (anonymous):

ok, I got 10.01 :)

OpenStudy (ipwnbunnies):

Yay.

OpenStudy (anonymous):

Is this right? NO2 is an empirical formula?

OpenStudy (ipwnbunnies):

It depends on what the full question is lol.

OpenStudy (anonymous):

Which formula is an empirical formula? A. C2H6 B.C4H10 C.NO2 D.NO

OpenStudy (ipwnbunnies):

Hmm, I don't like this question. It is C or D.

OpenStudy (anonymous):

I think I am guessing C

OpenStudy (ipwnbunnies):

Ok, just go with your gut. I can be wrong on this.

OpenStudy (anonymous):

Then answer D it is

OpenStudy (ipwnbunnies):

I say it's a bad question because both of them can be empirical formulas.

OpenStudy (anonymous):

Number 5 out of 10.... A compound contains 27.3% carbon and 72.7% oxygen by mass. What is the empirical formula of the compound? I got CO when I did it.

OpenStudy (ipwnbunnies):

I got CO2

OpenStudy (ipwnbunnies):

Assume we have 100 grams of the compound. 27.3 g of it would be C, and 72.7 g of it would be O. When we convert these masses to moles: C: 2.275 mol && O: 4.54 mol

OpenStudy (ipwnbunnies):

Our fake formula looks like this: C2.275 O4.54; Divide by the smallest subscript, so 2.275 in this case. C1O2 --> CO2

OpenStudy (anonymous):

That makes perfect sense to me, Thanks:)

OpenStudy (anonymous):

Your helping me lots

OpenStudy (ipwnbunnies):

Yeah, it makes perfect sense when I show the answer lol.

OpenStudy (ipwnbunnies):

I know, at least you're understanding.

OpenStudy (anonymous):

Yes for sure, Next one.. Given the reaction: 2H2(g) + O2 (g) = 2 H2O(_) What is the total number of liters of H2 (g) at STP needed to produce 6.02X10^23 molecules of H2O(_)? When I did this problem myself I got 22.4L

OpenStudy (ipwnbunnies):

I believe that is correct.

OpenStudy (ipwnbunnies):

No, hold on.

OpenStudy (ipwnbunnies):

Ok, yes that's correct lol. Thought we didn't have a mole of H2O for a second.

OpenStudy (anonymous):

Yayy!! What is the percent composition by mass of Carbon in NH4HCO3 (gram-formula mass = 79 grams/mole)? When I did this problem I got the answer 15.2%

OpenStudy (ipwnbunnies):

Yay

OpenStudy (anonymous):

Im getting better at doing these:)

OpenStudy (anonymous):

Only 3 left.. Given the reaction 2 C8H18(g)+2509g0= 16CO2(g) + 18H2O(g) What volume of C8H18(g) will completely react to produce exactly 18 liters of H2O(g)? When I tried it I got... 2.0L

OpenStudy (ipwnbunnies):

That's + 25 O2, right? lol

OpenStudy (ipwnbunnies):

Is this at STP, or it didn't say?

OpenStudy (ipwnbunnies):

I don't think it matters, it should be constant pressure. 2L is right

OpenStudy (anonymous):

No its 250 down 2(g)

OpenStudy (anonymous):

and if that doesn't change anything then 2L is my answer

OpenStudy (ipwnbunnies):

What? It has to be 25 O2. I know this reaction lol. What do you mean 250 down 2?

OpenStudy (ipwnbunnies):

It shouldn't change the answer, so yeah.

OpenStudy (ipwnbunnies):

It's 25 O2, probably written as 2502 accidently.

OpenStudy (anonymous):

Thanks ill email my teacher about that... Phosphoric acid reacts with sodium hydroxide according to the equation ; H3PO4 + 3NaOH = Na3PO4 = 3H2O. if 2.00 mol H3PO4 is made to react with 3.00mol NaOH, identify the limiting reactant. When I tried it I got NaOH

OpenStudy (ipwnbunnies):

Yep

OpenStudy (anonymous):

Ahhh... im happy this is all clicking in my head Last one.... If 60.0g FeCO3 is heated with an excess of oxygen, 45.0g of Fe2O3 is produced. If 52.5grams were expected, what is the percent yield of this reaction? My answer is 91.8%

OpenStudy (ipwnbunnies):

Err no, and this is the easiest question yet lol. Percent yield of the product = actual Fe2O3/expected Fe2O3 * 100%

OpenStudy (anonymous):

im confused because the actual divided by the expected times 100 = 100

OpenStudy (ipwnbunnies):

No, 45g/52.5g * 100% =

OpenStudy (anonymous):

Thank you, Thank you and Thank you for all the help... Ill be studying these types of problems so im ready for my exam on Thursday

OpenStudy (ipwnbunnies):

No prob